40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Jan 13 23:30:53 2020, Yehonathan, Summary, PSL, Changes made since Gautam left 
    Reply  Tue Jan 14 20:17:09 2020, gautam, Summary, General, IFO recovery DCerrors.png
       Reply  Wed Jan 15 10:04:19 2020, gautam, Summary, General, POX / POY locking restored singleArms.png
Message ID: 15119     Entry time: Mon Jan 13 23:30:53 2020     Reply to this: 15121
Author: Yehonathan 
Type: Summary 
Category: PSL 
Subject: Changes made since Gautam left 

As per Gautam's request, I list the changes that were made since he left:

1. The AOM driver was connected to a signal generator.

2. The first order beam from the AOM was coupled into the PMC while the zero-order beam is blocked. We might want to keep this configuration if the pointing stability is adequate.

3. c1psl got Burt restored to Dec 1st.

4. Megatron got updated.

Currently, c1susaux seems unresponsive and needs to be rebooted.

ELOG V3.1.3-